OpenStudy (anonymous):

Find a formula in terms of x for the function w(x) = p(p(x)), where p(x) = 2x + 1.

OpenStudy (jusaquikie):

i think you just plug in the terms so w(x) = [2(2x+1)+1]

OpenStudy (jusaquikie):

then if you need to simplyfy and solve for x that's not that hard either, (2(2x+1)+1) (4x+2+1) 4x+3 4x=-3 x=-3/4

OpenStudy (anonymous):

I see where you plug in the 2x+1 for p(x) but how do you decide what the first p in the equation is, confused on that p

OpenStudy (jusaquikie):

p(x) is the function then it is ran through itself again

OpenStudy (jusaquikie):

think of v=p(x) p(v) is the same as the p(p(x)) in the problem given

OpenStudy (anonymous):

awesome, thank you, I see what your talking about and I can understand it, you are a stress reducer and life saver!!!

OpenStudy (jusaquikie):

p, we all had a tough time with functions at one point, later they will seem so natural that it's like addition

OpenStudy (anonymous):

I'm gonna take your word for it and keep in mind that maybe they will get easier

OpenStudy (jusaquikie):

it all comes down to putting a number in and getting a number out what your working on now is putting a number in then getting a number out and then putting that number back in to get out the answer